OpenStudy (anonymous):

CAN SOMEONE HELP ME!? I GIVE MEDALS!!!!! Provide three unique examples of polynomials of varying degree and classification. State the classification and degree of each. (ie. 3x3+2y2x2 classification: binomial, degree: fourth)

OpenStudy (anonymous):

Degree 3 3x^3 + 2x^2 + x - 2= 0 Degree 2 3y^2 + 2y + 2 = 0 Degree 1 x + 2 = 0

OpenStudy (aum):

Example Degree Classification x^3 - 2x 3 Binomial 7x^2 -8x + 5 2 Trinomial 15x 1 Monomial

sammixboo (sammixboo):

Nice job, guys.

OpenStudy (anonymous):

THANK YOU :)